OpenStudy (anonymous):

find the two unknown lengths. A right triangle has one leg that is twice as long as the other leg. The hypotenuse is 2 square root 5 inches.

hero (hero):

Hint: \[x^2 + (2x)^2 = (2\sqrt{5})^2\]

OpenStudy (anonymous):

68?

hero (hero):

\[x^2 + 4x^2 = 20\]\[5x^2 = 20\]\[x^2 = 20/5\]\[x^2 = 4\]\[x = \sqrt{4}\]\[x = 2\] Therefore short leg = 2 long leg = 4

OpenStudy (anonymous):

Thanks Hero .

OpenStudy (anonymous):

What about for: A right triangle has a hypotenuse that is 3 feet longer than one leg. The other leg is 4 feet.

hero (hero):

The procedure is the same. Use pythagorean theorem and go from there.

OpenStudy (anonymous):

Thanks .

hero (hero):

Are you going to try it? If so, I'd like to see what you come up with for a solution.

OpenStudy (anonymous):

I'm gonna try it .

OpenStudy (anonymous):

Wait, what do i do first?

hero (hero):

1. Write down Pythagorean Theorem: \[a^2 + b^2 = c^2\]

hero (hero):

I suppose you want the second step instead.

OpenStudy (anonymous):

3^2+4^2=c^2?

hero (hero):

\[4^2 + x^2 = (x + 3)^2\]

hero (hero):

Try your best to make sense of that and then let me know what the value of x is.

OpenStudy (anonymous):

(x+3)^2 becomes x^2+9 right?

hero (hero):

Hint: (x+3)^2 = (x+3)(x+3)

OpenStudy (anonymous):

x^2+9? lol

hero (hero):

(x+3)^2 = (x+3)(x+3) = x(x+3)+3(x+3) = x^2 + 3x + 3x + 9 = x^2 + 6x + 9

hero (hero):

That would be the Distribution/FOIL Method

OpenStudy (anonymous):

I see... I've been doing math all day and now it's all confusing now.

hero (hero):

I'm sorry to hear that. =[

OpenStudy (anonymous):

Yeah, this is why i can't figure this out right now .

hero (hero):

So are you telling me that you never learned how to FOIL?

OpenStudy (anonymous):

Yes, I have learned FOIL .

hero (hero):

Well, those steps I did above are FOIL. Maybe slightly different but only because I used a step by step procedure which involves distributive rule.

OpenStudy (anonymous):

I even tried the foil and that's what i got too.

hero (hero):

You tried FOIL and got x^2 + 9 ?

OpenStudy (anonymous):

Actually i got x^2+6x+9

hero (hero):

Oh, okay. Perfect! Now you can continue solving the problem =]

OpenStudy (anonymous):

By doing what? lol

OpenStudy (anonymous):

Oh wait i think i got this

OpenStudy (anonymous):

1.5?

hero (hero):

So after expanding the right side we have: \[4^2 + x^2 = x^2 + 6x + 9\] Now, just use basic algebra methods to solve for x. Hint: Put like terms all on one side.

OpenStudy (anonymous):

7/6?

hero (hero):

Yes

OpenStudy (anonymous):

So 6 & 7/6?

hero (hero):

You'd be better off rounding 7/6 to a decimal.

OpenStudy (anonymous):

6 & 1.16?

hero (hero):

Where do you get six from?

hero (hero):

I suppose you never understood how I came up with (x+3) to begin with.

OpenStudy (anonymous):

Still confused....